# A randomized lattice rule without component-by-component construction

### v1

@@ 1. Introduction
There have been some recent works addressing this issue. In [5], Dick et al. introduced a randomized CBC construction to randomly choose a good generating vector $\boldsymbol{z}$ for the case $\alpha>1/2$. This is an online algorithm, where each sample $\boldsymbol{z}$ is constructed with a computational cost of $O(dN\log N)$ by the fast Fourier transform approach from [21] for a randomly chosen $N$. In [16], Kuo et al. eliminated the randomness in choosing a generating vector and demonstrated that, with a carefully constructed single generating vector $\boldsymbol{z}$, a randomized lattice rule with a randomly chosen prime $N\in\{\lceil M/2\rceil+1,\ldots,M\}$, which is the only source of randomness in the algorithm, still achieves nearly the optimal rate of the randomized error. The corresponding generating vector can be constructed offline, while the CBC algorithm requires a computational cost of $O(dM^{4}/\log M)$. This result is further extended to the RMSE for all $\alpha>0$ by applying a random shift in [22]. In passing, a randomized trapezoidal rule with random $N$ to approximate integrals with respect to one-dimensional Gaussian measure has been investigated in [10].
In this paper, we revisit the property that $|\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}|\geq\eta(N-1)^{d}$, indicating that, with $\eta=1/2$, at least half of the possible choices for the generating vector $\boldsymbol{z}$ from $\{1,\ldots,N-1\}^{d}$ are already good. In fact, the set $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ is defined as a collection of the generating vectors whose corresponding lattice rules achieve nearly the optimal order of the deterministic worst-case error in $H_{\alpha,\boldsymbol{\gamma}}$. This property was previously shown in [8, Theorem 2] and has been utilized in a recent work by the author and L’Ecuyer [11], where the median of several independently and randomly chosen rank-1 lattice rules was shown able to adjust to the smoothness and weight parameters of $H_{\alpha,\boldsymbol{\gamma}}$ without any need to specify them. In this paper, we exploit the same property in the following way to introduce yet another randomized lattice rule: by independently and randomly drawing generating vectors $r$ times from the set $\{1,\ldots,N-1\}^{d}$ and then comparing the worst-case errors of rank-1 lattice rules with these generating vectors in $H_{\alpha,\boldsymbol{\gamma}}$, the generating vector that minimizes the worst-case error among them belongs to the set $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ with a very high probability $1-(1-\eta)^{r}$. Although there is an exponentially small failure probability, its contribution to the overall error bound can be mitigated through averaging over randomness with an appropriate choice of $r$.
@@ 2.1. Weighted Korobov space
For a real number $\alpha>1/2$ and a sequence of non-negative real numbers $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},\ldots)\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$, where each $\gamma_{j}$ is assumed to be bounded above by $1$ for normalization, let $r_{\alpha,\boldsymbol{\gamma}}:\mathbb{Z}^{d}\to\mathbb{R}_{\geq 0}$ be defined by
@@
where we set the empty product to $1$ and $|k_{j}|^{\alpha}/\gamma_{j}=\infty$ for any $k_{j}\in\mathbb{Z}$ if $\gamma_{j}=0$. Then the weighted Korobov space $H_{\alpha,\boldsymbol{\gamma}}$ is defined as a reproducing kernel Hilbert space with the inner product
@@
Note that even for $20\leq\alpha\leq 1/$ 2, the weighted Korobov space can be defined as above, and it holds that $H_{\alpha,\boldsymbol{\gamma}}\subseteq L_{2}([0,1]^{d})$ with equality for $\alpha=0$ and $\gamma_{j}>0$ for all $j\in\mathbb{N}$. However, the Fourier series of $f\in H_{\alpha,\boldsymbol{\gamma}}$ does not necessarily converge absolutely, and the equality in (2) should be understood to hold almost everywhere. As we focus on the case $\alpha>1/2$ throughout this paper, the Fourier series of $f\in H_{\alpha,\boldsymbol{\gamma}}$ always converges absolutely.
@@ 2.2. Rank-1 lattice rule
$B_{N,\eta,\alpha,\boldsymbol{\gamma},\lambda}:=\left(\frac{1}{(1-\eta)(N-1)}\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}\left(2\zeta(\alpha/\lambda)\right)^{|u|}\right)^{\lambda},$
@@
In this proposition, the bound $B_{N,\eta,\alpha,\boldsymbol{\gamma},\lambda}$ on the worst-case error approaches the rate $O(N^{-\alpha})$ arbitrarily closely as a function of $N$ when $\lambda\to\alpha-0$. Moreover, it can be shown that this bound is further bounded independently of the dimension $d$ for all $1/2\leq\lambda<\alpha$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$.
@@ 3.1. Algorithm and basic properties
    $\mathcal{P}_{M}:=\left\{\ \lceil M/2\rceil<p\leq M\,\colon\,\text{$p$ is prime}\right\}.$
@@
    $\boldsymbol{z}^{*}:=\arg\min_{\boldsymbol{z}_{j},1\leq j\leq r}e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}},P_{N,\boldsymbol{z}_{j}}).$
Our new, online randomized rank-1 lattice rule (without shift), denoted simply by $A^{\mathrm{ran}}_{r,M}$, is defined by a family of $Q_{P_{N,\boldsymbol{z}^{*}}}$, where $N$ and $\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{r}$ are random variables and $\boldsymbol{z}^{*}$ is determined as described in the third step of Algorithm 3.1 with inputs $\alpha,\boldsymbol{\gamma},r,M$. With an additional random shift, in which case our new randomized lattice rule is denoted by $\tilde{A}^{\mathrm{ran}}_{r,M}$, we draw $\boldsymbol{\Delta}=(\Delta_{1},\ldots,\Delta_{d})$ randomly from the uniform distribution over $[0,1]^{d}$ and $Q_{P_{N,\boldsymbol{z}^{*}}}$ is replaced by $Q_{P_{N,\boldsymbol{z}^{*}}+\boldsymbol{\Delta}}$, where
@@
for any $10<\eta<$ 1 .
@@
Recall that the condition $\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}$ is equivalent to $\boldsymbol{k}\cdot\boldsymbol{z}_{1}\equiv 0\pmod{N}$, which holds for all $\boldsymbol{z}_{1}\in\{1,\ldots,N-1\}^{d}$ if $\boldsymbol{k}\equiv 0\pmod{N}$, i.e., if $k_{j}$ is a multiple of $N$ for all $j$. Otherwise, if $\boldsymbol{k}\not\equiv 0\pmod{N}$, i.e., if there exists at least one index $\ell$ such that $k_{\ell}$ is not any multiple of $N$, then the condition $\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}$ is further equivalent to $k_{\ell}z_{1,\ell}\equiv-\boldsymbol{k}_{-\{\ell\}}\cdot\boldsymbol{z}_{1,-\{\ell\}}\pmod{N}$, which has at most one solution $\tilde{z}_{1,\ell}\in\{1,\dots,N-1\}$ for any $\boldsymbol{z}_{1,-\{\ell\}}\in\{1,\ldots,N-1\}^{d-1}$ since $N$ is a prime. This implies that we have
$\mathbb{P}_{\boldsymbol{z}_{1}\mid N}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}\right]\leq\begin{cases}1&\text{if $\boldsymbol{k}\equiv 0\ (\mathrm{mod}\ N)$,}\\ (N-1)^{-1}&\text{otherwise.}\end{cases}$
@@
$\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]$ $\leq 1-\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\left(1-\frac{1}{N-1}\right)^{r}$
$=\frac{1}{|\mathcal{P}_{M}|}\left[\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}1+\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\left(1-\left(1-\frac{1}{N-1}\right)^{r}\right)\right]$
$\leq\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}1+\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}\frac{r}{N-1}$
$\leq\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv 0\ (\mathrm{mod}\ N)\end{subarray}}1+\frac{r}{\lceil M/2\rceil}.$
@@ 3.2. Error bounds
(3) $(1-\eta)M\geq 2\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}\left(2\zeta(\alpha/\lambda)\right)^{|u|}$
for all $1/2\leq\lambda<\alpha$. Then it holds that
$e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ $\leq\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}(2\zeta(\alpha/\lambda))^{|u|}\right)^{\lambda-\delta}$
$\qquad+(1-\eta)^{r}\left(\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{2}(2\zeta(2\alpha))^{|u|}\right)^{1/2},$
@@
Let us focus on the first term in (4). By the definition of $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ and Lemma 2.4, it holds for any given $N\in\mathcal{P}_{M}$ and $\boldsymbol{z}^{*}\in\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ that
@@
for all $1/2\leq\lambda<\alpha$. This means that, under the condition $\boldsymbol{z}^{*}\in\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$, we have $\boldsymbol{k}\not\in P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}$ if there exists $\lambda\in[1/2,\alpha)$ such that $1/r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})>B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}$ holds. We note here that the condition (3) ensures that $B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\leq 1.$ Together with this observation, by applying Lemma 3.3 and Cauchy–Schwartz inequality, we obtain
@@
$\leq\mathbb{E}\left[\sum_{\begin{subarray}{c}P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}|\hat{f}(\boldsymbol{k})|\right]$
$=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]|\hat{f}(\boldsymbol{k})|$
$\leq c\frac{r}{M}\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}|\hat{f}(\boldsymbol{k})|\log(1+\|\boldsymbol{k}\|_{\infty})$
$\leq c\frac{r}{M}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}(|\hat{f}(\boldsymbol{k})|r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}\right)^{1/2}$
×$\qquad\qquad\times\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}\right)^{1/2}$
$\leq c\frac{r}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}\right)^{1/2},$
for any $1/2\leq\lambda<\alpha$. Although we omit the remaining argument for bounding the last sum over $\boldsymbol{k}$ from above since it follows exactly the same way as those in [5, 14], see also [6, Chapter 11], we can show that, for any $1/2<\lambda<\alpha$ and $0<\delta<\min(\lambda-1/2,1)$, there exists a constant $c^{\prime}_{\lambda,\delta}>0$, depending only on $\lambda$ and $\delta$, such that
$\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\qquad\qquad\leq c^{\prime}_{\lambda,\delta}\left(B_{\lceil M/2\rceil+1,\eta,\alpha,\boldsymbol{\gamma},\lambda}\right)^{2(\lambda-\delta)-1}\sum_{\emptyset\neq u\subseteq\{1{:}d\}}\gamma_{u}^{1/\lambda}(2\zeta(\alpha/\lambda))^{|u|}$
holds. This gives an upper bound on the first term in (4) as

### v3

@@ 1. Introduction
There have been some recent works addressing this issue. In [5], Dick et al. introduced a randomized CBC construction to randomly choose a good generating vector $\boldsymbol{z}$ for the case $\alpha>1/2$. This is an online algorithm, where each sample $\boldsymbol{z}$ needs to be constructed at runtime with a computational cost of $O(dN\log N)$ by the fast Fourier transform approach from [21] for a randomly chosen $N$. In [16], Kuo et al. eliminated the randomness in choosing a generating vector and demonstrated that, with a carefully constructed single generating vector $\boldsymbol{z}$, a randomized lattice rule with a randomly chosen prime $N\in\{\lceil M/2\rceil+1,\ldots,M\}$, which is the only source of randomness in the algorithm, still achieves nearly the optimal rate of the randomized error. The corresponding generating vector can be constructed offline ahead of time, while the CBC algorithm requires a computational cost of $O(dM^{4}/\log M)$. This result is further extended to the RMSE for all $\alpha>0$ by applying a random shift in [22]. In passing, a randomized trapezoidal rule with random $N$ to approximate integrals with respect to one-dimensional Gaussian measure has been investigated in [10].
In this paper, we revisit the property that $|\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}|\geq\eta(N-1)^{d}$, indicating that, with $\eta=1/2$, at least half of the possible choices for the generating vector $\boldsymbol{z}$ from $\{1,\ldots,N-1\}^{d}$ are already good. In fact, the set $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ is defined as a collection of the generating vectors whose corresponding lattice rules achieve nearly the optimal order of the deterministic worst-case error in $H_{\alpha,\boldsymbol{\gamma}}$. This property was previously shown in [8, Theorem 2] and has been utilized in a recent work by the author and L’Ecuyer [11], where the median of several independently and randomly chosen rank-1 lattice rules was shown to be able to adjust to the smoothness and weight parameters of $H_{\alpha,\boldsymbol{\gamma}}$ without any need to specify them. In this paper, we exploit the same property in the following way to introduce yet another randomized lattice rule: by independently and randomly drawing generating vectors $r$ times from the set $\{1,\ldots,N-1\}^{d}$ and then comparing the worst-case errors of rank-1 lattice rules with these generating vectors in $H_{\alpha,\boldsymbol{\gamma}}$, the generating vector that minimizes the worst-case error among them belongs to the set $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ with a very high probability $1-(1-\eta)^{r}$. Although there is an exponentially small failure probability, its contribution to the overall error bound can be mitigated through averaging over randomness with an appropriate choice of $r$.
@@ 2.1. Weighted Korobov space
For a real number $\alpha>1/2$ and a sequence of non-negative real numbers $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},\ldots)\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$, let $r_{\alpha,\boldsymbol{\gamma}}:\mathbb{Z}^{d}\to\mathbb{R}_{\geq 0}$ be defined by
@@
where we set the empty product to $1$ and $|k_{j}|^{\alpha}/\gamma_{j}=\infty$ for any $k_{j}\in\mathbb{Z}$ if $\gamma_{j}=0$. Later in this paper, more precisely in our theoretical analysis on randomized error criteria in Section 3.2, we assume that each $\gamma_{j}$ is bounded above by $1$. This ensures that $r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1$ for all $\boldsymbol{k}$.
Then the weighted Korobov space $H_{\alpha,\boldsymbol{\gamma}}$ is defined as a reproducing kernel Hilbert space with the inner product
@@
Note that even for $20\leq\alpha\leq 1/$ 2, the weighted Korobov space can be defined as above, and it holds that $H_{\alpha,\boldsymbol{\gamma}}\subseteq L_{2}([0,1]^{d})$ with equality for $\alpha=0$ and $\gamma_{j}>0$ for all $j\in\mathbb{N}$. For $20\leq\alpha\leq 1/$ 2, the functions in $H_{\alpha,\boldsymbol{\gamma}}$ are not necessarily periodic, unlike the case for $\alpha>1/2$ where we assume periodicity. Moreover, the Fourier series of $f\in H_{\alpha,\boldsymbol{\gamma}}$ does not necessarily converge absolutely if $20\leq\alpha\leq 1/$ 2, and the equality in (2) should be understood to hold almost everywhere. As we focus on the case $\alpha>1/2$ throughout this paper, the Fourier series of $f\in H_{\alpha,\boldsymbol{\gamma}}$ always converges absolutely.
@@ 2.2. Rank-1 lattice rule
$B_{N,\eta,\alpha,\boldsymbol{\gamma},\lambda}:=\left(\frac{1}{(1-\eta)(N-1)}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\right)^{\lambda},$
@@
In this proposition, the bound $B_{N,\eta,\alpha,\boldsymbol{\gamma},\lambda}$ on the worst-case error approaches the rate $O(N^{-\alpha})$ arbitrarily closely as a function of $N$ when $\lambda\to\alpha^{-}$. Moreover, it can be shown that this bound is further bounded independently of the dimension $d$ for all $1/2\leq\lambda<\alpha$ if $\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}<\infty$, since then
$\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\leq\mathrm{e}^{2\zeta(\alpha/\lambda)\sum_{j=1}^{d}\gamma_{j}^{1/\lambda}}\leq\mathrm{e}^{2\zeta(\alpha/\lambda)\sum_{j=1}^{\infty}\gamma_{j}^{1/\alpha}},$
but, unfortunately, $\zeta(\alpha/\lambda)\to\infty$ if $\lambda\to\alpha^{-}$.
@@ 3.1. Algorithm and basic properties
    $\mathcal{P}_{M}:=\left\{p\in\mathbb{N}_{p}\,\colon\,\lceil M/2\rceil<p\leq M\right\},$
    where $\mathbb{N}_{p}$ denotes the set of prime numbers.
@@
    $\boldsymbol{z}^{*}:=\arg\min_{1\leq j\leq r}e^{\mathrm{wor}}(H_{\alpha,\boldsymbol{\gamma}};P_{N,\boldsymbol{z}_{j}}).$
Our new, online randomized rank-1 lattice rule (without shift), denoted simply by $A^{\mathrm{ran}}_{r,M}$, is defined as a family of $Q_{P_{N,\boldsymbol{z}^{*}}}$, where $N$ and $\boldsymbol{z}^{*}$ are random variables. Here, $N$ is always a prime number, as it is drawn from the set $\mathcal{P}_{M}$, consisting of primes in the range $\lceil M/2\rceil<p\leq M$. “Online” refers to the fact that the selection of the generating vectors needs to be performed at runtime, rather than being predetermined beforehand. Specifically, the realizations of $N$ and $\boldsymbol{z}^{*}$ are generated as described in Algorithm 3.1 with inputs $\alpha,\boldsymbol{\gamma},r,M$. With an additional random shift, in which case our new randomized lattice rule is denoted by $\tilde{A}^{\mathrm{ran}}_{r,M}$, we draw $\boldsymbol{\Delta}=(\Delta_{1},\ldots,\Delta_{d})$ randomly from the uniform distribution over $[0,1]^{d}$ and $Q_{P_{N,\boldsymbol{z}^{*}}}$ is replaced by $Q_{P_{N,\boldsymbol{z}^{*}}+\boldsymbol{\Delta}}$, where
@@
for any $10<\eta<$ 1, where $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ is defined in Proposition 2.5.
@@
Recall that the condition $\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}$ is equivalent to $\boldsymbol{k}\cdot\boldsymbol{z}_{1}\equiv 0\pmod{N}$, which holds for all $\boldsymbol{z}_{1}\in\{1,\ldots,N-1\}^{d}$ if $\boldsymbol{k}\equiv\boldsymbol{0}\pmod{N}$, i.e., if $k_{j}$ is a multiple of $N$ for all $j$. Otherwise, if $\boldsymbol{k}\not\equiv\boldsymbol{0}\pmod{N}$, i.e., if there exists at least one index $\ell$ such that $k_{\ell}$ is not any multiple of $N$, then the condition $\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}$ is further equivalent to $k_{\ell}z_{1,\ell}\equiv-\boldsymbol{k}_{-\{\ell\}}\cdot\boldsymbol{z}_{1,-\{\ell\}}\pmod{N}$, which has at most one solution $\tilde{z}_{1,\ell}\in\{1,\dots,N-1\}$ for any $\boldsymbol{z}_{1,-\{\ell\}}\in\{1,\ldots,N-1\}^{d-1}$ since $N$ is a prime. This implies that we have
$\mathbb{P}_{\boldsymbol{z}_{1}\mid N}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}_{1}}^{\perp}\right]\leq\begin{cases}1&\text{if $\boldsymbol{k}\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)$,}\\ (N-1)^{-1}&\text{otherwise.}\end{cases}$
@@
$\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]$ $\leq 1-\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}\left(1-\frac{1}{N-1}\right)^{r}$
$=\frac{1}{|\mathcal{P}_{M}|}\left[\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}1+\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}\left(1-\left(1-\frac{1}{N-1}\right)^{r}\right)\right]$
$\leq\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}1+\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\not\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}\frac{r}{N-1}$
$\leq\frac{1}{|\mathcal{P}_{M}|}\sum_{\begin{subarray}{c}N\in\mathcal{P}_{M}\\ \boldsymbol{k}\equiv\boldsymbol{0}\ (\mathrm{mod}\ N)\end{subarray}}1+\frac{r}{\lceil M/2\rceil}.$
@@ 3.2. Error bounds
[…]
Then it holds that
$e^{\mathrm{ran}}(H_{\alpha,\boldsymbol{\gamma}};A^{\mathrm{ran}}_{r,M})$ $\leq\frac{rC_{\lambda,\delta}}{(1-\eta)^{\lambda-\delta-1/2}M^{\lambda-\delta+1/2}}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)^{\lambda-\delta}$
$\qquad+(1-\eta)^{r}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{2}\zeta(2\alpha)\right)^{1/2},$
@@
Let us focus on the first term in (4). By the definition of $\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$, and by Lemma 2.4 and Proposition 2.5, it holds for any given $N\in\mathcal{P}_{M}$ and $\boldsymbol{z}^{*}\in\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$ that
@@
for all $1/2\leq\lambda<\alpha$. By defining
$H_{M,\eta,\alpha,\boldsymbol{\gamma}}:=\inf_{1/2\leq\lambda<\alpha}\left(\frac{2}{(1-\eta)M}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)\right)^{\lambda},$
this means that, under the condition $\boldsymbol{z}^{*}\in\mathcal{Z}_{N,\eta,\alpha,\boldsymbol{\gamma}}$, we have $\boldsymbol{k}\not\in P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}$ if $1/r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})>H_{M,\eta,\alpha,\boldsymbol{\gamma}}$ holds. We note here that the condition (3) ensures that $H_{M,\eta,\alpha,\boldsymbol{\gamma}}\leq 1.$ Together with this observation, by applying Lemma 3.3 and Cauchy–Schwarz inequality, we obtain
@@
$\leq\mathbb{E}\left[\sum_{\begin{subarray}{c}\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}|\hat{f}(\boldsymbol{k})|\right]$
$=\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\mathbb{P}\left[\boldsymbol{k}\in P_{N,\boldsymbol{z}^{*}}^{\perp}\right]|\hat{f}(\boldsymbol{k})|$
$\leq c\frac{r}{M}\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}|\hat{f}(\boldsymbol{k})|\log(1+\|\boldsymbol{k}\|_{\infty})$
$\leq c\frac{r}{M}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}(|\hat{f}(\boldsymbol{k})|r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}\right)^{1/2}$
×$\qquad\qquad\times\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}\right)^{1/2}$
$\leq c\frac{r}{M}\|f\|_{\alpha,\boldsymbol{\gamma}}\left(\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}\right)^{1/2}.$
Although we omit the remaining argument for bounding the last sum over $\boldsymbol{k}$ from above since it follows exactly the same way as those in [5, 14], see also [6, Chapter 11], we can show that, for any $1/2<\lambda<\alpha$ and $0<\delta<\min(\lambda-1/2,1)$, there exists a constant $c^{\prime}_{\lambda,\delta}>0$, depending only on $\lambda$ and $\delta$, such that
$\sum_{\begin{subarray}{c}\boldsymbol{k}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}\\ r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})\geq 1/H_{M,\eta,\alpha,\boldsymbol{\gamma}}\end{subarray}}\frac{(\log(1+\|\boldsymbol{k}\|_{\infty}))^{2}}{(r_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k}))^{2}}$
$\qquad\qquad\leq c^{\prime}_{\lambda,\delta}\left(H_{M,\eta,\alpha,\boldsymbol{\gamma}}\right)^{2-(2\delta+1)/\lambda}\prod_{j=1}^{d}\left(1+2\gamma_{j}^{1/\lambda}\zeta(\alpha/\lambda)\right)$
holds. We note that the assumption $\gamma_{j}\leq 1$ for all $j$ is implicitly used here. This gives an upper bound on the first term in (4) as